### v2

# Persistent Homology-Based Indicator of Orientational Ordering in Two-Dimensional Quasi-Particle Systems applied to Skyrmion Lattices

Abstract
Two-dimensional (2D) quasi-particle systems, such as magnetic skyrmions, exhibit topological phase transitions between unique 2D phases. However, to identify the phases, an easy and computationally low-cost methodology for capturing the configurational properties of lattice ordering and constructing such an appropriate descriptor that can be easily calculated is, though, not obvious. Here, we propose an indicator for topological phase transitions using persistent homology (PH). PH offers novel insights beyond conventional indicators by capturing topological features derived from the configurational properties of the lattice. The proposed persistent-homology-based topological indicator, which selectively counts stable features in a persistence diagram, effectively traces the lattice’s topological changes, as confirmed by comparisons with the conventionally used measure of the ordering (the magnitude of the orientational order parameter $\langle|\Psi_{6}|\rangle$), typically used to identify lattice phases. We demonstrate the applicability of our indicator to experimental data, showing that it yields results consistent with those of simulations. This experimental validation highlights the robustness of the proposed method for real physical systems beyond idealized simulated systems. While our method is demonstrated in the context of skyrmion lattice systems, the approach is general and can be extended to other two-dimensional systems composed of interacting quasi-particles. As a key advantage, our indicator offers lower computational complexity than the conventionally used measures.
1  Introduction
Magnetic skyrmions are chiral, vortex-like spin textures characterized by topologically enhanced stability [1, 2], often treated as quasi-particles due to their particle-like behavior [3, 4]. These structures have garnered significant attention from researchers for their energy-efficient manipulation via spin-transfer and spin–orbit torques [5], offering promising pathways toward next-generation information storage and unconventional computing devices [6]. A possible application of skyrmions is the racetrack memory as a high-density memory technology [7, 8, 9]. Such future skyrmionic applications are based on the behavior of multiple skyrmions. It is, therefore, important to investigate their physical properties and their collective arrangements to achieve such technological applications using skyrmions.
From a fundamental physics perspective, ensembles of skyrmions, known as skyrmion lattices, can exhibit complex phase behavior, including two-dimensional melting transitions [10]. These transitions are theoretically captured by the Kosterlitz–Thouless–Halperin–Nelson–Young
(KTHNY) framework, which describes the emergence of intermediate hexatic phases and the role of topological defects such as dislocations and disclinations [10]. Although this behavior has been extensively studied in colloidal systems, skyrmion lattices offer distinct advantages due to their tunability and controllable dynamics that can be fully captured in real time and real space, making them an ideal platform for investigating two-dimensional phase transitions [11].
Despite intensive research, a computationally easy methodology for describing the configurational properties of lattices remains elusive. Conventional methods (e.g., using the local orientational order parameter $\psi_{6}(r)$ [5, 11, 12, 13]) are limited by their reliance on ensemble-averaged quantities and are computationally expensive. That is because they normally average the quantities encoded to each particle over all the particles to identify the state in the systems.
To address this, we introduce a framework based on Topological Data Analysis (TDA), a concept from algebraic topology used to analyze the geometric structure of objects, specifically applying Persistent Homology (PH) to extract configurational properties of lattice configurations [14, 15, 16, 17, 18, 19]. The PH provides a multi-scale view of topological features by analyzing how connected components and loops emerge and disappear across a filtration. This approach has seen successful applications in biological and materials systems [20, 21, 22, 23, 24]. PH can capture and describe the microscopic processes involved in the phase transitions [16]. In this work, we adopt a particle-based approach, modeling skyrmions as interacting quasi-particles and focusing on their positional configuration that does not require knowledge of the full spin texture. This abstraction enables an efficient computational treatment and highlights the essential geometric features responsible for phase behavior [3, 4, 25]. To extract physically meaningful insights from the persistence diagrams, we propose a topological indicator [26, 27], the Persistent Generator Count with Relative Stability (PGCRS), which selectively counts only the robust topological features in each homology dimension. PGCRS reliably detect phase transitions in the lattice, correlates strongly with conventional measures, and offers a significantly reduced computational complexity. A distinctive aspect of our approach is inverse analysis [28], to trace persistent generators back to specific real-space configurations, enabling a direct physical interpretation of the microscopic structures responsible for topological phase behavior. This conceptual shift from averaging local order to counting persistent topological features offers a new perspective for understanding 2D phase transitions. In addition to numerical simulations, we apply our framework to experimentally acquired skyrmion configurations, demonstrating the practical applicability of our method for real physical systems. While the focus of this work is on skyrmion systems, the underlying method can be extended to other two-dimensional quasi-particle systems.
This paper is organized as follows: we first describe the simulation setup, experimental procedure and analysis framework. We then present detailed results, including the persistence diagrams, inverse structural analysis, and PGCRS, and conclude with a discussion of implications and potential extensions.
2  Methods
In this work, we treat skyrmions as quasi-particles where the effective inter-particle interactions reflect those of real skyrmion systems, without considering their detailed spin textures [3, 25, 29]. This abstraction allows us to focus on the configurational and topological properties relevant to phase transitions, while significantly simplifying the computational analysis.
The skyrmion lattice states used for persistent homology analysis are generated via Brownian dynamics simulation, specifically a Thiele model-based approach, using the HOOMD-blue package [30]. In this work, the Magnus term is assumed negligible under the high-density skyrmion condition, where the resulting short mean free path limits its influence, and is therefore omitted [31]. Consequently, the Thiele model simulation effectively reduces to a standard Brownian dynamics simulation in the overdamped limit [31]. This simulation models a two-dimensional system of quasi-particles interacting under conditions that reproduce a range of ordering according to different phases, including liquid, hexatic, and solid states [10, 11].
In addition to the simulation data, we apply our method to experimental skyrmion configurations, allowing for a direct comparison with the numerically simulated configurations in the persistent homology analysis.
In the subsequent sections, we first describe the simulation setup, including the inter-particle potential and system initialization, as well as the experimental procedure. We then introduce the local orientational order parameter, its correlation function, and the averaged absolute value of the local orientational order parameter. Finally, we describe the persistent homology analysis and present the construction of the Persistent Generator Count with Relative Stability, our proposed topological indicator.
2.1  Brownian Dynamics Simulation
The Brownian dynamics simulation is performed by using the HOOMD-blue package [30]. The modelled system is a 2D system of $N=65000$ particles with hexagonal packing under periodic boundary conditions. The governing model is classical Brownian dynamics, in which $\beta\varepsilon=1$ is set, where $\beta=1/k_{B}T=1$ denotes the inverse thermal energy and $\varepsilon$ represents the interaction strength [29]. The interaction potential in this system is described by
$U(r)=\varepsilon({\sigma}/{r})^{n},$ (1)
where $\sigma$ represents the characteristic interaction range, $n$ corresponds to the steepness of the potential, and $r$ is the distance from the referenced particle [3, 5, 29]. In this work, $\sigma=1.0$, and $n=8.0$ are chosen. It is worth noting that two commonly used effective skyrmion-skyrmion interaction models are the power-law decaying function, used in our simulations, and the exponentially decaying function [3, 5, 11]. The simulation time step is chosen as $dt=10^{-4}$. The well-equilibrated initial state is prepared under an equilibration process for a duration of $10^{7}$ time steps to ensure relaxation to a steady state. After equilibration, snapshots are recorded every $10^{5}$ time steps for additional $10^{7}$ time steps, ensuring no correlations in the state between snapshots. To generate the state spanning a wide range of topological phases, 22 density values are scanned from $\phi$ = 1.1000 to 1.7000. The system is initialized with the 65000 quasi-particles under constant thermal activation over different systems in which the topological phases change depending on the value of density, which is defined as the number of particles relative to the size of the simulation box (specifically, the density value is tuned by the size of the simulation box).
2.2  Experimental Data
The experimental skyrmion configuration data used in this work are obtained from magnetic multilayer stacks: Ta(5)/Co20Fe60B20(0.9)/Ta(0.07)/MgO(2)/Ta(5) [3, 11]. Magnetic fields are applied both in-plane (IP) and out-of-plane (OOP) using an Evico Magnetics GmbH Kerr microscope, with the sample maintained at a constant temperature of 333.5 K. Skyrmions are nucleated by applying an IP field pulse under a constant OOP field. The lattice is subsequently equilibrated using an oscillating OOP field at 100 Hz (with amplitudes up to $60\,\upmu\mathrm{T}$), combined with a constant OOP offset field. The skyrmion size is precisely controlled via the applied OOP magnetic field. The magnetic field conditions are varied every 62.5 s (corresponding to 1000 frames) to gather sufficient statistics for analyzing the topological phases. The positions of individual skyrmions are identified using a machine-learning-based, pixel-wise classification algorithm [32]. The detailed analysis procedure can be found in [11]. The experimentally acquired coordinates are analyzed using the same framework as that employed for the simulation data, enabling a direct comparison between experimental and numerical datasets.
2.3  Local Orientational Order Parameter $\psi_{6}(r)$, Correlation Function $c_{6}(r)$, Averaged Absolute Value of Local Orientational Order Parameter $\langle|\Psi_{6}|\rangle$
To identify the topological phases, including liquid, hexatic, and solid phases, the following quantity, termed orientational correlation function, is typically used [5, 11, 12, 13],
$c_{6}(r=\left|r_{k}-r_{j}\right|)=\langle\psi_{6}(r_{k})\psi_{6}^{*}(r_{j})\rangle,$ (2)
where
$\psi_{6}(r_{j})=\frac{1}{n}\sum_{m=1}^{n}\exp(i6\theta_{m}^{j}),$ (3)
n is the number of nearest neighbors of the referenced particle $j$, and $\theta_{m}^{j}$ is the angle between $r_{m}-r_{j}$ and a fixed arbitrary axis. Here, the nearest neighbors are determined using Voronoi tessellation [31]. The factor 6 in the exponent reflects the six-fold rotational symmetry of a perfect hexagonal lattice. When a particle’s six nearest neighbors form an ideal hexagonal arrangement, an absolute value of $\psi_{6}$ of 1 is obtained. The KTHNY theory explains the typical behavior of a correlation function for each topological phase, specifically liquid, hexatic, and solid phases exhibit an exponential decay, algebraic decay, and constant behavior close to 1 as a function of distance $r$, respectively [10, 13].
In addition to identifying the topological phase via the orientational correlation function, another measure of the ordering is the averaged absolute value of the local orientational order parameter $\langle|\Psi_{6}|\rangle$, as described in the following equation [5].
$\langle|\Psi_{6}|\rangle=\left|\frac{1}{N}\sum_{j=1}^{N}\psi_{6}(r_{j})\right|,$ (4)
where $N$ is the number of particles. This measure of the ordering can be used as a quick indication of the topological phase, as the computational cost is lower compared to the full correlation function [5]. As for other reasons for using it, $\langle|\Psi_{6}|\rangle$ can be used as a single scalar value, making it easier to compare across different conditions, such as different densities, and it does not require additional fitting, unlike the correlation function. Regarding the fitting to the correlation function, the extracted decay value may not be very precise, especially if noise or finite-size effects are present.
2.4  Alpha Filtration of the Multi-Degree Homology in Persistent Homology Analysis for Two-Dimensional Coordinates of Quasi-Particles
In the persistent homology analysis, an alpha-filtration of the 0th- and 1st-degree homology is often applied to the two-dimensional coordinates of quasi-particles. From this, the corresponding persistence diagrams (PD0 and PD1) can be generated. The inverse analysis method of persistent homology is applied to the selected frames of our data, which can identify the original structure corresponding to a specific generator in the persistence diagram.
2.5  Construction of Persistent Homology-Based Indicator
There has been some work [16, 26, 27] investigating the relationship between the phase transition and configuration space topology. In some classes of systems, phase transitions are accompanied by topology changes in the configuration manifold. To investigate the topological behavior of systems, we use some topological quantities defined in the configurational space. We refer the reader to Refs.[26, 27] for a detailed description of the following equations that we use:
Euler characteristics (EC) are integer-valued topological invariants, described as follows [26, 27]:
$\chi(X):=\sum_{k}(-1)^{k}\beta_{k}(X),$ (5)
where $X$ is a topological space, and $\beta_{k}$ is $k$-th Betti number, which counts the number of topological features. For instance, $\beta_{0}$, $\beta_{1}$, and $\beta_{2}$ correspond to the number of connected components, one-dimensional loops, and two-dimensional loops, respectively. In this paper, only the connected components and one-dimensional loops, thus $\beta_{0}$, and $\beta_{1}$ are considered, as the quasi-particle system is two-dimensional. The EC can be expressed as the alternating sum of the $k$-th Betti number. One of the important properties of the EC is that it is a topological invariant. For example, if two different states have a different value of the EC $\chi_{X_{1}}$, and $\chi_{X_{2}}$, the two states are topologically non-equivalent [27]. Therefore, by monitoring the evolution of the EC, the topology changes within the configurational space can be detected.
In a filtration $\left\{X_{t}\right\}$ of the persistent homology, the EC curve is defined as
$\chi(t):=\chi(X_{t})=\sum_{k}{(-1)^{k}\beta_{k}(X_{t})},$ (6)
which tracks the behavior of the EC in time (filtration parameter) [26]. To obtain the integer-valued topological invariants in the framework of the persistent homology, we define the Persistent Generator Count with Relative Stability, as described in the following equation
$\tilde{\chi}_{\text{rel}}(X_{\text{PH}}):=\sum_{k}{(-1)^{k}\beta_{k}^{\text{rel}}(X_{\text{PH}})},$ (7)
where $\beta_{k}^{\text{rel}}(X_{\text{PH}})$ counts the number of generators in the persistence diagram (PD) that satisfy specific stability conditions, as described in the supplementary section. While PGCRS is inspired by Betti number formalism, further theoretical justification is needed to fully interpret its implications in the context of topological phase transitions.
2.6  Gaussian Process Regression
The first derivatives of the PGCRS $\langle\tilde{\chi}\rangle$ and the conventionally used measure of lattice ordering $\langle|\Psi_{6}|\rangle$ are estimated from a smoothed predictive curve constructed using Gaussian Process Regression (GPR) with a Radial Basis Function (RBF) kernel. For the simulation data, three different RBF kernels are employed to account for the non-uniform sampling intervals along the density axis $\phi$. In contrast, the experimental dataset, which consists of more sparsely sampled points, is modeled using a single RBF kernel. The expressions for the predictive mean curve and the associated uncertainty estimates from the GPR model are provided in Ref.[33].
3  Results and Discussion
We now apply a new approach to analyze skyrmion ensembles. This section is organized as follows: we first examine the behavior of the conventionally used order parameter. Next, we analyze persistent diagrams and investigate the role of generators in the persistence diagram (PD) through inverse analysis. Finally, we compare the Persistent Generator Count with Relative Stability (PGCRS) to the conventional measure and validate the results using experimental data.
3.1  Conventional Used Measure of Lattice Ordering in Simulation Data
Fig. 1 shows the correlation function of the local orientational order parameter as a function of the distance for three skyrmion density values ($\phi=1.1000$, $1.1960$, and $1.7000$) and the averaged absolute value of the local orientational order parameter as a function of density values. Each state in the left panel corresponds to a distinct phase, as indicated in Ref.[29]. The correlation functions for the three density values show the exponential decay, algebraic decay, and constant value close to 1, respectively, as determined by the dotted black line representing $r^{-\frac{1}{4}}$ [11]. According to the KTHNY theory prediction [10, 11, 13, 29], these behaviors correspond to the liquid, hexatic, and solid phases, respectively, with the decay rate exceeding the critical $r^{-\frac{1}{4}}$ threshold indicating the liquid phase. The right panel shows the averaged absolute value of the local orientational order parameter $\langle|\Psi_{6}|\rangle$ as a function of density value. It indicates that $\langle|\Psi_{6}|\rangle$ increases, which indicates more ordered configurational alignment, as the density values increase.
Figure 1: The correlation function of the local orientational order parameter as a function of the distance for the three skyrmion density values $\phi=1.1000$, $1.1960$, and $1.7000$ and the averaged absolute value of the local orientational order parameter as a function of density values. The three states correspond to the liquid, hexatic, and solid phases, respectively, as determined by the dotted black line, which represents a power law decay with a critical exponent $r^{-\frac{1}{4}}$ [11]. The left panel shows that $\langle|\Psi_{6}|\rangle$ increases, which indicates more ordered configurational alignment, as the density values increase.
3.2  Persistent Diagrams of 0th- and 1st-Degree Homology in Simulation Data
Fig. 2 displays the persistence diagrams (PDs) of the 0th- and 1st- degree homology, respectively, generated from the three skyrmion density values $\phi=1.1000$, $1.1960$, and $1.7000$, corresponding to the liquid, hexatic, and solid phases. The ”Birth” and ”Death” represent the specific time in a filtration process of the Persistent Homology (PH) in which the topological features emerge and disappear, respectively, and the color map represents the multiplicity of the generators (scatter plots) in the PD. The magnitude of the values of ”Birth” and ”Death” corresponds to the length of the disk located at the coordinates of the quasi-particles, which increases during a filtration process of the PH. It is worth noting that, in this work, the coordinates of the system have been normalized to eliminate a possible effect from the absolute density values. It is observed that the distribution of generators in the PD of 0th-degree homology (PD0) for the disordered structure (e.g., the states at skyrmion density values $\phi=1.1000$, and $1.1960$) is broader compared with the ordered structure (e.g., the state at density value $\phi=1.7000$). It is also confirmed that the disordered structures appear to have more generators in PD of 1st-degree homology (PD1) with a large lifetime, which is a distance from the diagonal line in PD. In contrast, the ordered structure likely has fewer generators with a long lifetime in PD1.
[…]
3.3  Connection of Generators on PD with Microscopic Configurations in Simulation Data
Fig. 3 presents the inverse analysis [28] for two skyrmion density values $\phi=1.1000$ and $1.7000$. This analysis traces back the specific generators in PD to the original structure in the real-space configuration. In each state, the points labeled (a), (b), and (c) correspond to persistent homology generators with large lifetime (a), small lifetime and small birth value (b), and small lifetime and large birth value (c), respectively. It is clearly observed that the generators with a large lifetime originate from a complex structure, and the generators with a small lifetime originate from a simple structure. The result confirms the previous subsection, as the disordered structure tends to a complex alignment, while the ordered structure likely leads to a perfect alignment.
Figure 3: Inverse analysis for the two skyrmion density values $\phi=1.1000$ and $1.7000$, tracing specific generators in PD back to real-space configurations. In each state, the points labeled (a), (b), and (c) correspond to persistent homology generators with large lifetime (a), small lifetime and small birth value (b), and small lifetime and large birth value (c), respectively.
3.4  Persistent Homology-Based Indicator and Comparison with Conventional Used Measure of Lattice Ordering in Simulation Data
The comparison between the conventional method for phase identification and persistent homology is performed quantitatively by using the Persistent Generator Count with Relative Stability, as discussed next:
Fig. 4 presents the correlation between the Persistent Generator Count with Relative Stability and conventionally used measure of the ordering $\langle|\Psi_{6}|\rangle$ and the first derivatives of both indicators with the Pearson correlation coefficient. The details of the first derivatives and the Gaussian Process Regression, used to estimate the first derivatives, are provided in the supplementary section. The figure shows a positive proportional relationship, as indicated by the Pearson correlation coefficient $r=0.981$ and $r=0.998$. The high sensitivity of the first derivatives supports the validity of the persistent homology-based approach, as phase transitions are characterized by abrupt, non-analytic changes in system properties [34].
Figure 4: Relationship between the Persistent Generator Count with Relative Stability $\langle\tilde{\chi}\rangle$ and the conventionally used measure of the ordering $\langle|\Psi_{6}|\rangle$. The relationship between the first derivatives of both indicators is also shown using Gaussian Process Regression. The Pearson correlation coefficient $r=0.981$ and $r=0.998$ indicate a good correlation between both values.
3.5  Validation through Application to Experimental Skyrmion Lattices
[…]
In each state, the points labeled (a), (b), and (c) correspond to persistent homology generators with a large lifetime (a), small lifetime and small birth value (b), and small lifetime and large birth value (c), respectively. As consistent with those of the simulation data, it is also observed that the generators with a large lifetime originate from a complex structure, and the generators with a small lifetime originate from a simple structure. The result confirms the previous subsection, as the disordered structure tends to be complex alignment, while the ordered structure likely leads to a perfect alignment.
[…]
Fig. 7 presents the Persistent Generator Count with Relative Stability as a function of $\langle|\Psi_{6}|\rangle$, along with the first derivatives of both indicators. A good positive correlation is observed, as confirmed by the Pearson correlation coefficients $r=0.993$ and $r=0.861$ for the values and their derivatives, respectively. The details of the Gaussian Process Regression, used to estimate the first derivatives, are described in the supplementary section. These results highlight the consistency between the persistent homology-based indicator and the conventional orientational order parameter. As previously discussed in the simulation data, the high sensitivity, in particular the first derivatives, supports the validity of the persistent homology-based approach, as phase transitions are characterized by abrupt, non-analytic changes in system properties [34].
Figure 7: Correlation between the Persistent Generator Count with Relative Stability $\langle\tilde{\chi}\rangle$ and the conventional orientational order parameter $\langle|\Psi_{6}|\rangle$ in the experimental data. First derivatives are compared using Gaussian Process Regression. A good correlation is indicated by the Pearson correlation coefficients $r=0.993$ and $r=0.861$.
Finally, we analyze the computational cost of our indicator compared to the conventionally used indicators. Table 1 shows the comparison in computational complexity for the persistent homology-based indicator and the conventinally used $\langle|\Psi_{6}|\rangle$ [12, 35]. It is obvious that the newly constructed indicator has achieved a significant reduction in computational complexity, as the dimension of the system in this work is 2, and therefore the cost is $O(N)$. Furthermore, in the simulation data, the actual runtime for the persistent homology-based indicator is 2.34 seconds per frame, whereas the conventional used one requires $4.49\times 10^{3}$ seconds in our computational environment.
The ability of persistent homology analysis to measure the ordering may originate from its shared geometric foundation with the orientational order parameter, as both rely on Voronoi tessellation, complemented by techniques from algebraic topology [36]. One limitation, however, is that the persistent homology-based indicator may require comparisons across different system states to provide meaningful insights. The conventionally used measure of the ordering $\langle|\Psi_{6}|\rangle$ have a clear physical interpretation due to its normalization, which ensures values between 0 (disordered state) and 1 (perfect hexagonal order) [5].
Table 1: Comparison of computational complexity for the persistent homology-based indicator and the averaged absolute value of the local orientational order parameter  [12, 35]. Here, NN is the number of particles in the system, and $d$ is the dimension. In this work (simulation data), $N=65000$ and $d=2$.
4  Conclusion
In this work, we propose a topological indicator, a Persistent Generator Count with Relative Stability (PGCRS), for characterizing phases and phase transitions in two-dimensional quasi-particle systems. As a model system, we use skyrmions, which can be well described as quasi-particles that form lattices in 2D systems. By modeling skyrmions as interacting quasi-particles and applying persistent homology (PH) to their spatial configurations, PGCRS selectively counting stable topological features, providing an interpretable, noise-resistant, and computationally efficient measure of lattice ordering. It correlates strongly with the conventional orientational order parameter $\langle|\Psi_{6}|\rangle$, and reliably traces phase transitions across solid, hexatic, and liquid states. Our inverse analysis reveals that persistent generators with long lifetimes correspond to disordered, complex configurations, while short-lived features are associated with regular, ordered structures. The applicability of our approach to experimental data is demonstrated using experimental skyrmion lattices, confirming that the PGCRS produces consistent and reliable phase characterization under realistic experimental conditions.
While this work focuses on skyrmion lattice systems, the methodology is broadly applicable to other two-dimensional systems composed of repulsively interacting quasi-particles. Future extensions could incorporate insights, such as the Euler characteristic, to deepen the connection between structural ordering and thermodynamic quantities [26, 27].
In summary, PGCRS provides a practical and theoretically grounded framework for efficient, interpretable, and broadly applicable phase analysis in quasi-particle systems.